# Physical Processes Behind the Co-Evolution of Halos, Galaxies and Supermassive Black Holes in the IllustrisTNG Simulation

### v2

@@ Abstract
We explore the co-evolution of dark matter halos, their central galaxies, and central supermassive black holes (SMBHs) using the IllustrisTNG (TNG) simulation. We find that the evolutionary histories of individual galaxies in the $M_{\rm BH}$-$M_{*}$ plane can be decomposed into four distinct phases, separated by three transition points. We identify the driving processes of galaxy evolution within each phase and derive the conditions necessary and sufficient for transitions to subsequent phases. The first phase is dominated by star formation, with its duration primarily determined by the mass of the SMBH seed and the surrounding gas environment. The second phase is characterized by rapid SMBH growth, and the transition to the next phase occurs when the thermal-mode feedback of active galactic nucleus (AGN) can unbind gas from the galaxy. The third phase involves self-regulation of the SMBH, and the transition to the quenched phase occurs when the kinetic-mode feedback of AGN counterbalances gas cooling within the subhalo. The final phase is dominated by mergers. We investigate the use of scaling relations among different mass components and evolutionary phases to understand processes implemented in TNG and other simulations, and discuss how current and forthcoming observations can be used to constrain models.
@@ 1 Introduction
Structure formation in the $\Lambda$CDM cosmology is predicted to follow a hierarchical paradigm (e.g. Peebles, 1980; Blumenthal et al., 1984; Mo et al., 2010), in which structures grow from smaller to larger scales, forming virialized dark matter halos of larger masses. Gas associated with dark matter, which may be heated by shocks associated with gravitational collapse, can cool radiatively and contract until it becomes self-gravitating and fragments to form stars. Under extreme conditions that are most likely fulfilled in the central regions of galaxies, the gas and/or stars can be turned into black holes (BHs); the latter can continue to grow via accretion of gas and/or mergers with stars and other BHs. This may lead to the formation of supermassive black holes (SMBHs) observed in the centers of many galaxies (e.g. Tonry, 1984, 1987; Magorrian et al., 1998; Ferrarese & Merritt, 2000; Gebhardt et al., 2000; Shankar, 2009; Kormendy & Richstone, 1995; Kormendy & Ho, 2013; Graham & Sahu, 2023a; Maiolino et al., 2024). In this scenario of structure formation, small and compact objects in the cosmic hierarchy are embedded in the environments defined by the more extended and less dense structures. Inversely, energetic processes on small scales can affect their environments through feedback processes, forming a bidirectional connection in the hierarchy. Examples of such feedback include the AGN feedback from an accreting SMBH, which can push/heat the gas within its host galaxy (e.g. Fabian, 2012; Liu et al., 2013; Heckman & Best, 2014; Kubo et al., 2022; Concas et al., 2022; Donahue & Voit, 2022) and halo (e.g. Li et al., 2022; Luo et al., 2024; Liu et al., 2024), and the feedback from star formation, which can change the composition and state of the interstellar and circumgalactic medium (e.g. Fielding et al., 2017; Hopkins et al., 2020; Li et al., 2020; Deng et al., 2024) and even modify the distribution of dark matter (e.g. El-Zant et al., 2001; Mo & Mao, 2004; Agertz & Kravtsov, 2016; Acevedo et al., 2023; Yang et al., 2024). The formation and evolution of SMBHs, galaxies and halos are thus indispensable to one another, and have to be modeled together in order to understand the physical processes behind observational phenomena.
@@
Numerous theoretical efforts have been made to model structure formation in the current cosmological framework. Numerical (e.g. Springel & Hernquist, 2003; Springel et al., 2005; Springel, 2010; Schaye et al., 2015; Volonteri et al., 2016; Anglés-Alcázar et al., 2017; Weinberger et al., 2017; Hopkins et al., 2018; Pillepich et al., 2018a; Weinberger et al., 2018; Davé et al., 2019; McAlpine et al., 2018; Zinger et al., 2020; Blank et al., 2019; Stern et al., 2021; Li et al., 2022; Yang et al., 2023; Schaye et al., 2023; Luo et al., 2024; Bhowmick et al., 2024), semi-analytical (e.g. Kauffmann et al., 1993; Kang et al., 2005; Croton et al., 2006; Somerville et al., 2008; Guo et al., 2011; Benson, 2012; Lu et al., 2015a; Hirschmann et al., 2016; Lacey et al., 2016; Henriques et al., 2015; Lacey et al., 2016; Mo et al., 2024; Chen et al., 2024a, b; Lucia et al., 2024), analytical (e.g. Mo et al., 1998; Hernquist & Springel, 2003; Davé et al., 2012; Sharma & Theuns, 2020; Fukugita & Kawasaki, 2022; Bower et al., 2017; Dekel et al., 2023) and empirical (e.g. Jing et al., 1998; Zheng et al., 2005; Lu et al., 2014, 2015b; Hearin et al., 2016; Moster et al., 2018; Xu et al., 2018; Behroozi et al., 2019) models have been developed to describe the physical processes on different scales and to discover, test and verify the assumptions and prescriptions governing these processes. Various observations of SMBHs, galaxies and halos have been used to constrain these models. Examples of observations include the spatially resolved properties of individual galaxies and clusters (e.g. Brough et al., 2007; Li et al., 2015; Zhou et al., 2019; Lin et al., 2019; Mirakhor & Walker, 2020; Fahrion et al., 2020; Zhou et al., 2021; Churazov et al., 2021; Hong et al., 2023; Lin et al., 2023; Li et al., 2023; Harris, 2023; Zhang et al., 2023; Jing & Li, 2024); the spatial number density of galaxies binned by different quantities, such as the galaxy luminosity or stellar mass function, and AGN luminosity or SMBH mass function (e.g. Shankar et al., 2008; Li & White, 2009; Kelly & Merloni, 2012; Bongiorno et al., 2016; Tucci & Volonteri, 2017; Chen et al., 2019; Yang et al., 2021; Sicilia et al., 2022; Padmanabhan & Loeb, 2023; Meng et al., 2023; Wang et al., 2024a); the spatial correlation of galaxies and AGNs (e.g. Li et al., 2006; Kirkpatrick et al., 2017; Wang & Li, 2019; Mountrichas, 2023; García-Bernete et al., 2024; Saxena et al., 2024); the joint distribution or scaling relations between different quantities, such as the SMBH mass-stellar mass relation (e.g. Kormendy & Ho, 2013; Graham & Sahu, 2023a, b; Hong et al., 2023), the SMBH mass-halo mass relation (e.g. Ferrarese, 2002; Gaspari et al., 2019; Marasco et al., 2021; Zhang et al., 2024; Li et al., 2024a), and the stellar mass-halo mass relation (e.g. Yang et al., 2003, 2007, 2008; Moster et al., 2010; Behroozi et al., 2013; Wechsler & Tinker, 2018; Behroozi et al., 2019; Zhang et al., 2021, 2022).
@@
Comparisons between observations and models have led to many important conclusions about the physical processes on different scales and the interaction between them. For example, the star formation efficiency, defined as the ratio of stellar mass to halo mass, is found observationally to be an unimodal function of halo mass, peaked at around a halo mass of $M_{\rm h,c}\approx 10^{12}M_{\odot}$ (e.g. Yang et al., 2003, 2012; Behroozi et al., 2013). The peak star formation efficiency is about a few percent, much lower than the value of $\approx 0.16$ allowed by the cosmic baryon fraction. Two different channels of feedback, one from stars (e.g. Kauffmann & Charlot, 1998; Hopkins et al., 2012, 2014; Anglés-Alcázar et al., 2017; Hopkins et al., 2021) and the other from AGNs (e.g. Bower et al., 2017; Weinberger et al., 2018; Davé et al., 2019) are thus proposed to suppress the star formation, and are inferred to be dominant at low- and high-masses, respectively. Another example is the relation between SMBH mass and galaxy bulge/total stellar mass, which is observed to roughly follow a power law (e.g. Graham, 2012; Scott & Graham, 2013; Graham & Scott, 2013; Kormendy & Ho, 2013; Graham & Sahu, 2023a, b; Hong et al., 2023). Different mechanisms, such as star formation and SMBH growth regulated by AGN feedback (e.g. Di Matteo et al., 2005; Weinberger et al., 2018; Hong et al., 2023), and the galaxy-galaxy and SMBH-SMBH mergers (e.g. Peng, 2007), have been proposed to explain the relation, in particular the power-law index. Precise measurements of the relation over a wide range of redshift from observations are thus critical to distinguish theoretical models, although such measurements are difficult to make at high redshift (e.g. Schaye et al., 2010; Kormendy & Ho, 2013; Greene et al., 2020). The stellar mass-halo mass relation and the SMBH mass-stellar mass relation also imply a correlation in mass between SMBHs and halos, as indeed found observationally by using halo masses obtained from the abundance matching (Behroozi et al., 2019), satellite kinematics (Ferrarese, 2002), gravitational lensing (Zhang et al., 2024; Li et al., 2024a), and luminosity in X-ray (e.g. Gaspari et al., 2019; Voit et al., 2024), and SMBH mass from velocity dispersion (e.g. Zahid et al., 2016, 2018; Sohn et al., 2020; Shankar et al., 2020; Zhang et al., 2024). Physical pictures connecting baryon cycles from SMBH scales to halo scales have been suggested to explain these observations (e.g. Gaspari & Sadowski, 2017; Donahue & Voit, 2022).
James Webb Space Telescope (JWST) has opened a new era to study galaxy formation at high redshift. Over the past two years, observations with JWST have discovered a large number of galaxies that are much different from those seen in the local Universe. As galaxies at high redshift are featured by high densities and short dynamical time scales (e.g. Dekel et al., 2023), star formation and SMBH growth are expected to be more violent and bursty (Hopkins et al., 2023, 2024a), and the morphology of galaxies is expected to be more irregular (Ma et al., 2018). Indeed, studies based on JWST found that galaxies at high redshift appear to be more clumpy (e.g. Kuhn et al., 2024; Hirano & Yoshida, 2024; Bik et al., 2024), with some dominated by dense star-forming clumps (e.g. Vanzella et al., 2022; Lin et al., 2023; Mowla et al., 2024; Adamo et al., 2024). An abundant population of “little red dots”, featured by reddened spectra and small sizes, has been discovered at $z\gtrsim 4$, and has been inferred as AGNs and/or young stars obscured by dust (Matthee et al., 2024; Pérez-González et al., 2024; Li et al., 2024b). In addition, a recent analysis by Wang et al. (2024b) targeting at a young galaxy at $z\approx 8$ has suggested that its emission lines and spectral slope are most consistent with a cluster of Pop-III stars. The massive stars within this cluster are thus candidates for SMBH seeds. The above observations with JWST qualitatively support the theoretical expectations of galaxy formation in the $\Lambda$CDM paradigm, and, for the first time, join the pieces of information throughout the full growth histories of galaxies and SMBHs from their infancies to the present day. To use such observations to constrain theoretical models, it is crucial to identify processes that are responsible for salient observational properties, and to understand how different processes are linked together to produce the populations of galaxies and SMBHs observed at any given time.
@@ 2 Data and samples
A baryonic version of the SubLink algorithm is applied to TNG to build the subhalo merger trees (see Rodriguez-Gomez et al., 2015, for details). Each subhalo is linked with a set of progenitor subhalos, among which the one with the “most massive history” is defined as the first/main progenitor. To trace the evolution of a halo, its central galaxy, and the central SMBH within the central galaxy, we follow the main progenitor of the central subhalo recursively back in time. This extracts a branch of subhalos from the whole merger tree rooted in the subhalo in consideration, and we refer to it as the the main branch of this root subhalo/galaxy/SMBH.
@@
  • $M_{\rm h}$: halo mass, defined as the total mass enclosed in a sphere centered on the particle of minimal gravitational potential among all particles in the FoF halo, with a radius, $R_{\rm h}$, defined so that the mean density within the sphere is 200 times the mean density of the universe at the epoch when the halo is identified. The circular velocity of the halo at $R_{\rm h}$ is defined as the virial velocity $V_{\rm h}$.
@@
  • $M_{\rm BH}$: the sum of the masses of all black holes in a subhalo, used in this paper as the mass of the central SMBH. We have checked that only $<5$% galaxies in our sample host more than one black hole at any redshift in the main branch. Our results are thus not significantly changed if we use an alternative definition of SMBH mass.
@@
For the halo, stellar and SMBH masses, the definition of $\dot{M}(t_{n})$ yields the halo accretion rate (HAR), the star formation rate (SFR), and the black hole accretion rate (BHAR), respectively, and the definition of $\dot{M}(t_{n})/M(t_{n})$ yields their specific values: the specific halo accretion rate (sHAR), the specific black hole accretion rate (sBHAR), and the specific star formation rate (sSFR).
We define a galaxy to be quiescent/quenched if its sSFR is less than a threshold of $10^{-11}\,{\rm yr}^{-1}$, which is commonly used in literatures (e.g. Schaye et al., 2015; Cui et al., 2021; Habouzit et al., 2021; Bluck et al., 2024). We find that the star formation history produced by TNG can be fluctuating. Some galaxies, in particular the high-redshift ones, can have sudden drops in sSFR in some individual snapshots, thus making an excursion to the “quenched” state and returning to the “star-forming” status. To ensure that our conclusion is robust against such fluctuations, we adopt a stronger criterion for a galaxy to be quenched by requiring its sSFR to remain below the threshold value for at least three continuous snapshots.
@@
Figure 1: The main-branch histories of individual galaxies in the $M_{\rm BH}$-$M_{*}$ plane. For each branch, $M_{*}$ and $M_{\rm BH}$ are scaled by their values at a reference redshift when the ratio, $M_{\rm BH}/M_{*}$, reaches its minimum. Top panel shows the histories of randomly selected galaxies in TNG at $z=0$ with $M_{*}>10^{8.5}\,{\rm M_{\odot}}$, 10 for star-forming (blue) and 10 for quenched (red). Bottom panel shows the sample of quenched galaxies used in the analysis throughout this paper. Here, each red dot represents a galaxy at a snapshot in a main branch. The red dashed curve shows the median among all branches. See §2 for the details of sample and selection.
@@
Fig. 1 shows the main-branch histories of individual galaxies in the $M_{\rm BH}$-$M_{*}$ plane. The top panel plots the histories of galaxies randomly selected from all central galaxies with $M_{*}>10^{8.5}\,{\rm M_{\odot}}$ at $z=0$, 10 for star-forming (blue) and 10 for quenched (red). Here, for each history, $M_{*}$ and $M_{\rm BH}$ are scaled by their values at a reference redshift when $M_{\rm BH}/M_{*}$ reaches its minimum. For each history, the $M_{\rm BH}$-$M_{*}$ curve presents regular patterns that mark the different phases undergone by the galaxy. In the early time before the reference point, the curve is quite flat, indicating that the growth of SMBH is slower than that of the stellar component. A sharp transition appears at about the reference point, followed by a rapid growth of SMBH. In the later time, the difference between star-forming and quenched galaxies appears. For quenched galaxies, the growth of $M_{\rm BH}$ slows down, and eventually reaches a regime where significant “jumps” appear in both $M_{*}$ and $M_{\rm BH}$. For most star-forming galaxies, this later phase is missing.
@@
Much of the effort in this paper is to quantify the distinct phases, and identify the underlying mechanisms driving the evolution during each phase and the transition to the next. In this paper, the main branches to be analyzed are selected according to the following criteria:
@@
The requirement for the subhalo to be a central (criterion i) and to have never been a satellite (criterion v) is to avoid the effect of the environment which provides additional channels to quench the galaxy (see, e.g. Kauffmann et al., 2013; Ayromlou et al., 2023; Wang et al., 2023). The lower limit imposed on $M_{*}$ (criterion ii) ensures that the root subhalo has at least $\approx 1000$ stellar particles, thus alleviating the artifacts due to numerical resolution. The requirement of the complete history (criterion iv) and quenched state (criterion iii) ensures that none of the four phases (see §3) is missed in the history, which simplifies the analysis and presentation of this paper. The numbers of galaxies passing the above 5 criteria are 6291208, 8280, 1682, 1594 and 1429, respectively. In addition, we have excluded four galaxies that cannot be properly fitted by the function to be discussed in §3. The conclusions of this paper thus apply only to quenched galaxies at $z=0$, for which we have checked that the above selection criteria do not lead to significant bias. For star-forming galaxies, as shown in the top panel of Fig. 1, their histories appear similar to the quenched galaxies with some later phases truncated. Whether or not this similarity is quantitatively true, how the evolution of star-forming galaxies in each phase and at each transition is different from that of quenched galaxies, and how SMBH feedback acts differently between star-forming and quenched galaxies (e.g. Fabian, 2012; Dubois et al., 2015; Weinberger et al., 2018), remain to be open questions to be addressed further.
The lower panel of Fig. 1 shows the main-branch histories in the $M_{\rm BH}$-$M_{*}$ plane for all the galaxies selected in our sample. The dashed red curve shows the median among all the histories. We label the processes that drive the evolution of galaxies in different phases in this figure, and we will discuss them in detail in the following sections.

### v3

@@ Abstract
We explore the co-evolution of dark matter halos, their central galaxies, and central supermassive black holes (SMBHs) using the IllustrisTNG (TNG) simulation. We find that the evolutionary histories of individual galaxies in the $_{\rm BH}-MM_{*}$ plane can be decomposed into four distinct phases, separated by three transition points. We identify the driving processes of galaxy evolution within each phase and derive the conditions necessary and sufficient for transitions to subsequent phases. The first phase is dominated by star formation, with its duration primarily determined by the mass of the SMBH seed and the surrounding gas environment. The second phase is characterized by rapid SMBH growth, and the transition to the next phase occurs when the thermal-mode feedback of active galactic nucleus (AGN) can unbind gas from the galaxy. The third phase involves self-regulation of the SMBH, and the transition to the quenched phase occurs when the kinetic-mode feedback of AGN counterbalances gas cooling within the subhalo. The final phase is dominated by mergers. We investigate the use of scaling relations among different mass components and evolutionary phases to understand processes implemented in TNG and other simulations, and discuss how current and forthcoming observations can be used to constrain models.
@@ 1 Introduction
[…]
@@
Numerous theoretical efforts have been made to model structure formation in the current cosmological framework. Numerical (e.g. Schaye et al., 2015; Hopkins et al., 2018; Davé et al., 2019; Pillepich et al., 2018a; Blank et al., 2019), semi-analytical (e.g. White & Rees, 1978; Somerville et al., 2008; Henriques et al., 2015; Croton et al., 2016; Baugh et al., 2019), analytical (e.g. Mo et al., 1998; Hernquist & Springel, 2003; Sharma & Theuns, 2020; Bower et al., 2017; Dekel et al., 2023) and empirical (e.g. Zheng et al., 2005; Hearin et al., 2016; Xu et al., 2018; Mo et al., 2024; Chen et al., 2024a, b) models have been developed to describe the physical processes on different scales and to discover, test and verify the assumptions and prescriptions governing these processes. Various observations of SMBHs, galaxies and halos have been used to constrain these models. Examples of observations include the spatially resolved properties of individual galaxies and clusters (e.g. Lin et al., 2019; Fahrion et al., 2020; Jing & Li, 2024); the spatial number density of galaxies binned by different quantities, such as the galaxy luminosity or stellar mass function, and AGN luminosity or SMBH mass function (e.g. Chen et al., 2019; Padmanabhan & Loeb, 2023; Wang et al., 2024a); the spatial correlation of galaxies and AGNs (e.g. Li et al., 2006; García-Bernete et al., 2024; Saxena et al., 2024); the joint distribution or scaling relations between different quantities, such as the SMBH mass-stellar mass relation (e.g. Kormendy & Ho, 2013; Graham & Sahu, 2023a; Hong et al., 2023), the SMBH mass-halo mass relation (e.g. Zhang et al., 2024; Li et al., 2024a), and the stellar mass-halo mass relation (e.g. Yang et al., 2007, 2008; Moster et al., 2010; Wechsler & Tinker, 2018; Zhang et al., 2022).
@@
Comparisons between observations and models have led to many important conclusions about the physical processes on different scales and the interaction between them. For example, the star formation efficiency, defined as the ratio of stellar mass to halo mass, is found observationally to be an unimodal function of halo mass, peaked at around a halo mass of $M_{\rm h,c}\approx 10^{12}M_{\odot}$ (e.g. Yang et al., 2003, 2012; Behroozi et al., 2013). The peak star formation efficiency is about a few percent, much lower than the value of $\approx 0.16$ allowed by the cosmic baryon fraction. Two different channels of feedback, one from stars (e.g. Kauffmann & Charlot, 1998; Hopkins et al., 2012, 2014; Anglés-Alcázar et al., 2017; Hopkins et al., 2021) and the other from AGNs (e.g. Bower et al., 2017; Weinberger et al., 2018; Davé et al., 2019) are thus proposed to suppress the star formation, and are inferred to be dominant at low- and high-masses, respectively. Another example is the relation between SMBH mass and galaxy bulge/total stellar mass, which is observed to roughly follow a power law (e.g. Kormendy & Ho, 2013; Graham & Sahu, 2023a, b; Hong et al., 2023). Different mechanisms, such as star formation and SMBH growth regulated by AGN feedback (e.g. Di Matteo et al., 2005; Weinberger et al., 2018; Hong et al., 2023), and the galaxy-galaxy and SMBH-SMBH mergers (e.g. Peng, 2007), have been proposed to explain the relation, in particular the power-law index. Precise measurements of the relation over a wide range of redshift from observations are thus critical to distinguish theoretical models, although such measurements are difficult to make at high redshifts (e.g. Schaye et al., 2010; Kormendy & Ho, 2013; Greene et al., 2020). The stellar mass-halo mass relation and the SMBH mass-stellar mass relation also imply a correlation in mass between SMBHs and halos, as indeed found observationally by using halo masses obtained from the abundance matching (Behroozi et al., 2019), satellite kinematics (Ferrarese, 2002), gravitational lensing (Zhang et al., 2024; Li et al., 2024a), and luminosity in X-ray (e.g. Gaspari et al., 2019; Voit et al., 2024), and SMBH mass from velocity dispersion (e.g. Zahid et al., 2016, 2018; Sohn et al., 2020; Shankar et al., 2020). Although the correlation between SMBH mass and halo mass may be just a phenomenological coincidence, physical frameworks involving the synergy between halo potential and AGN feedback in regulating the baryon cycle from SMBH scales to halo scales have been proposed to interpret the correlation (e.g. Bower et al., 2017; Voit et al., 2024), and observational evidences supporting the interpretation have been identified (e.g. Bandara et al., 2009; Bogdán & Goulding, 2015; Zhang et al., 2024).
[…]
@@ 2 Data and samples
A baryonic version of the SubLink algorithm is applied to TNG to build the subhalo merger trees (see Rodriguez-Gomez et al., 2015, for details). Each subhalo is linked with a set of progenitor subhalos, among which the one with the “most massive history” is defined as the first/main progenitor. To trace the evolution of a halo, its central galaxy, and the central SMBH within the central galaxy, we follow the main progenitor of the central subhalo recursively back in time. This extracts a branch of subhalos from the whole merger tree rooted in the subhalo in consideration, and we refer to it as the main branch of this root subhalo/galaxy/SMBH.
@@
  • $\bm{M}_{\rm h}$: halo mass, defined as the total mass enclosed in a sphere centered on the particle of minimal gravitational potential among all particles in the FoF halo, with a radius, $R_{\rm h}$, defined so that the mean density within the sphere is 200 times the mean density of the universe at the epoch when the halo is identified. The circular velocity of the halo at $R_{\rm h}$ is defined as the virial velocity $V_{\rm h}$.
@@
  • $M_{\rm BH}$: the mass of the most massive black-hole particle in the subhalo.
@@
For the halo, stellar and SMBH masses, the definition of $\dot{M}(t_{n})$ yields the halo accretion rate (HAR), the star formation rate (SFR), and the black hole accretion rate (BHAR), respectively, and the definition of $\dot{M}(t_{n})/M(t_{n})$ yields their specific values: the specific halo accretion rate (sHAR), the specific black hole accretion rate (sBHAR), and the specific star formation rate (sSFR). Note that the integration of the growth rate defined here is ensured to be equal to the mass whose evolution pattern is the main focus of this paper.
To tell whether a galaxy is star-forming or quenched, we find its instantaneous star formation rate by summing over all gas cells within $2R_{*,1/2}$, and obtain the instantaneous sSFR by dividing the instantaneous SFR by $M_{*}$. We note that the instantaneous SFR is noisy due to the stochastic nature of the subgrid models, the finite resolution of the simulation and the defects of the structure finder algorithms that link the resolution elements. Some galaxies, in particular the high-redshift ones, can thus have sudden drops in sSFR in some individual snapshots, make an excursion to the “quenched” state and return to the “star-forming” status later. To mitigate the noise, we smooth the instantaneous sSFR along the main branch of the subhalo merger tree by a running Gaussian kernel with standard deviation equal to the dynamical timescale ($t_{\rm dyn,h}\equiv R_{\rm h}/V_{\rm h}$) of the host halo at each snapshot. We define a galaxy to be quiescent/quenched if the sSFR so obtained is less than the threshold of $10^{-11}\,{\rm yr}^{-1}$, which is commonly used in literatures (e.g. Schaye et al., 2015; Cui et al., 2021; Habouzit et al., 2021; Bluck et al., 2024). For each quenched galaxy, we identify its quenching time, $t_{\rm q}$, as the first time when its smoothed instantaneous sSFR falls below the threshold of $10^{-11}\,\,{\rm yr}^{-1}$.
@@
Figure 1: The main-branch histories of individual galaxies in the $_{\rm BH}-MM_{*}$ plane. For each branch, $M_{*}$ and $M_{\rm BH}$ are scaled by their values at a reference redshift when the ratio, $M_{\rm BH}/M_{*}$, reaches its minimum. Top panel shows the histories of randomly selected galaxies in TNG at $z=0$ with $M_{*}>10^{8.5}\,{\rm M_{\odot}}$, 10 for star-forming (blue) and 10 for quenched (red). Bottom panel shows the sample of quenched galaxies used in the analysis throughout this paper. Here, each red dot represents a galaxy at a snapshot in a main branch. The red dashed curve shows the median among all branches. See §2 for the details of sample and selection.
@@
Fig. 1 shows the main-branch histories of individual galaxies in the $_{\rm BH}-MM_{*}$ plane. The top panel plots the histories of galaxies randomly selected from all central galaxies with $M_{*}>10^{8.5}\,{\rm M_{\odot}}$ at $z=0$, 10 for star-forming (blue) and 10 for quenched (red). Here, for each history, $M_{*}$ and $M_{\rm BH}$ are scaled by their values at a reference redshift when $M_{\rm BH}/M_{*}$ reaches its minimum. For each history, the $_{\rm BH}-MM_{*}$ curve presents regular patterns that mark the different phases undergone by the galaxy. In the early time before the reference point, the curve is quite flat, indicating that the growth of SMBH is slower than that of the stellar component. A sharp transition appears at about the reference point, followed by a rapid growth of SMBH. In the later time, the difference between star-forming and quenched galaxies appears. For quenched galaxies, the growth of $M_{\rm BH}$ slows down, and eventually reaches a regime where significant “jumps” appear in both $M_{*}$ and $M_{\rm BH}$. For most star-forming galaxies, this later phase is missing.
@@
Much of the effort in this paper is to quantify the distinct phases and identify the underlying mechanisms driving the evolution during each phase and the transition to the next. In this paper, the main branches to be analyzed are selected according to the following criteria:
@@
The requirement for the subhalo to be a central (criterion i) and to have never been a satellite (criterion v) is to avoid the effect of the environment which provides additional channels to quench the galaxy (see, e.g. Kauffmann et al., 2013; Ayromlou et al., 2023; Wang et al., 2023). The lower limit imposed on $M_{*}$ (criterion ii) ensures that the root subhalo has at least $\approx 1000$ stellar particles, thus alleviating the artifacts due to numerical resolution. The requirement of the complete history (criterion iv) and quenched state (criterion iii) ensures that none of the four phases (see §3) is missed in the history, which simplifies the analysis and presentation of this paper. The numbers of galaxies passing the above 5 criteria are 6291208, 8280, 1682, 1594 and 1429, respectively. In addition, we have excluded two galaxies that cannot be properly fitted by the function to be discussed in §3. The conclusions of this paper thus apply only to quenched galaxies at $z=0$, for which we have checked that the above selection criteria do not lead to significant bias. For star-forming galaxies, as shown in the top panel of Fig. 1, their histories appear similar to the quenched galaxies with some later phases truncated. Whether or not this similarity is quantitatively true, how the evolution of star-forming galaxies in each phase and at each transition is different from that of quenched galaxies, and how SMBH feedback acts differently between star-forming and quenched galaxies (e.g. Fabian, 2012; Dubois et al., 2015; Weinberger et al., 2018), remain to be open questions to be addressed further.
The lower panel of Fig. 1 shows the main-branch histories in the $_{\rm BH}-MM_{*}$ plane for all the galaxies selected in our sample. The dashed red curve shows the median among all the histories. We label the processes that drive the evolution of galaxies in different phases in this figure, and we will discuss them in detail in the following sections.